Your lawn mower requires an oil and gas mixture to make it run correctly. The mixture requires 3 times as much gas as oil. Gas costs $1.43 and oil costs $.85 a gallon. How many full gallons of the mixture can be made for $20? This should be just a one variable equation... I have tried 1.43y + (1/3)(.85)y <=20 and I am not getting the correct answer?
Answer by stanbon(75887)

Your lawn mower requires an oil and gas mixture to make it run correctly.
The mixture requires 3 times as much gas as oil.
---
g = 3o
------
Gas costs $1.43 and oil costs $.85 a gallon.
Value: 1.43g + 0.85o <= 20
----
How many full gallons of the mixture can be made for $20?
-----
Substitute for "g" and solve for "o":
1.43(3o) + 0.85o <= 20
5.14o <= 20
o <= 3.89 gallons of oil
---
solve for "g"
(g/3) <= 3.89
g <= 11.67 gallons of gas
=========================
Cheers,
Stan H.
===========================